OpenStudy (lena772):

Simplify square root of 2 multiplied by the cube root of 2.

OpenStudy (anonymous):

\[\LARGE 2^{ \frac{ 1 }{ 2 } } *2^{\frac{ 1 }{ 3 }} = 2^{ \frac{ 1 }{ 6 }} =\sqrt[6]{2}\]

OpenStudy (the_fizicx99):

√2 * 3^√2 2^1/2 * 2^1/3 = 1 / 6 2^1/6 6√2^1 = 6√2

OpenStudy (anonymous):

oops actually that's incorrect

OpenStudy (anonymous):

let me redo that

OpenStudy (anonymous):

\[\LARGE 2^{ \frac{ 1 }{ 2 } } *2^{\frac{ 1 }{ 3 }} = 2^{ \frac{ 1 }{ 2 }+\frac{ 1 }{ 3 }} =2^{\frac{ 5 }{ 6 }}=\sqrt[6]{2^{5}}\]

OpenStudy (the_fizicx99):

The same base >.<

OpenStudy (anonymous):

you need to add the exponents, not multiply them, sorry for the confusion

OpenStudy (anonymous):

yes it's the same base, but that doens;t matter, look, \( \large 2^{3}*2^{4} =2^{7} \) right? not \( \large 2^{12} \)

OpenStudy (anonymous):

the same logic applies to roots